OpenStudy (anonymous):

What is the area of the sector of a circle if the radius is 16cm and the angle rotation is 150 degrees? I got this question wrong, so i am hoping for some help.

OpenStudy (anonymous):

i think its just the radius * how much it's rotated but the degrees should be radians 150 * (pi/180) = 5pi/6 so 16 * 5pi/6 is the answer

OpenStudy (anonymous):

which should come out to ~41

OpenStudy (anonymous):

That is what I had put on my homework and he counted it wrong. I think that is for length, not area.

OpenStudy (anonymous):

yes you're right one way you can do it though is to find the area of the whole circle, then find what 150/360 of it is. A = pi*r^2 A = pi*16^2 A = 301.44 but that is for 360 degrees, you want only 150. 150/360 = 5/12 so the real area is (5 * 301.44) / 12

OpenStudy (anonymous):

which should be roughly 125 square centimetres

OpenStudy (anonymous):

where did the five come from?

OpenStudy (anonymous):

5 comes from the ratio 5/12, which is just a simplified 150/360 as there are 360 total degrees in a circle, and you are looking for 150 of it (little less than half the circle) so you are looking for the area of 150/360 of a circle, or 5/12 you know the area of the full circle is 301.44 so (5/12) * 301.44 should give you the smaller area

OpenStudy (anonymous):

basically i set up a ratio: 301.44 sq cm for 360 degrees x sq cm for 150 degrees 301.44/360 = x/150 301.44*150/360 = x 301.44*(5/12) = x x being the area

OpenStudy (anonymous):

if that helps you picture it

OpenStudy (anonymous):

ok, I don't know what i did but i did not get 301.44

OpenStudy (anonymous):

301.44 is the area of the TOTAL circle you only want PART of the circle the part of the circle you are looking for, is 5/12 of the total circle. therefore, 5/12*301.44 is the answer which is about 125 square centimeters

OpenStudy (anonymous):

ok, thank you. I understand.